Solve the problem. The product of two consecutive integers is 5 more than their sum. Find the integers
Answer by Alan3354(69443)

The product of two consecutive integers is 5 more than their sum. Find the integers
------------
x*(x+1) = x + x+1 + 5
x^2 + x = 2x+6
x^2 - x - 6 = 0
(x-3)*(x+2) = 0
x = 3 --> 3 and 4
x = -2 --> -2 and -1